Two boats start together and race across a 60-km-wide lake and back. Boat A goes across at 60 km/h and returns at 60 km/h. Boat B goes across at 30 km/h, and its crew, realizing how far behind it is getting, returns at 90 km/h. Turnaround times are negligible, and the boat that completes the round trip first wins. 


1
Expert's answer
2022-01-10T09:09:25-0500

The boat, for which it takes the shortest time to pass the entire route, will win. The time for the boats:


"t_A=\dfrac {2d}{60}=\dfrac d{30}=0.033d.\\\space\\
t_B=\dfrac{d}{30}+\dfrac{d}{90}=\dfrac{2d}{45}=0.044d."


Boat B will lose.